# Security and Privacy on Generative Data in AIGC: A Survey

### v2

@@ V-B1 Generative Detection
Building a high-quality, diverse, and large-scale dataset is essential for training and evaluating generative data detectors. Although some datasets [80, 81] have been published, they have different limitations, e.g., targeting only a certain class of images or generators, and containing only a small amount of data. The Huawei team has publicly released GenImage [82], referred to as the “ImageNet dataset of the AIGC era”, which has significant advantages, including millions of images, richness of image content, and adoption of state-of-the-art generators. With the GenImage, researchers can effectively accelerate the advancement and assessment of advanced detectors in real-world scenarios. Generative image detection: The presence of artifacts in generative images is an important detection cue, which may derive from defects in the generation process or from a specific generative architecture. Corvi et al. [83] gave a preliminary trial to the problem of detecting generative images produced by DMs. Their study showed that the hidden artifact features of images of DMs are partially similar to those observed in images of GANs. In addition, the performance of detectors varies greatly in different scenarios, as these detectors focus on different cues and are hardly capable of strong generalization. Xi et al. [84] developed a robust dual-stream network consisting of a residual stream and a content stream to capture generic anomalies generated by AIGC. The residual stream utilizes the spatial rich model (SRM) to extract various texture information from images, while the content stream captures additional artifact traces at low frequencies, thus supplementing the residual stream with information that may have been missed. Sinitsa et al. [85] presented a rule-based method that can achieve high detection accuracy by training a small number of generative images (less than 512). The method employs the inductive bias of CNNs to extract fingerprints of different generators from the training set and applies it to detect generative images of the same model and its fine-tuned versions.
Figure 8: Generative data attributes to the training data from [86].
Analyzing distinctive features of generated images is also a viable approach to consider. Interestingly, Wang et al. [87] observed that the image generated by the DMs can be reconstructed by approximating the source model, while the real image cannot. Therefore, they proposed a novel image representation called diffusion reconstruction error (DIRE), which measures the distance between the input images and the reconstructed one. DIRE provides a reliable, simple, and generalized method to differentiate between real images and diffusion-generated images. Amoroso et al. [88] investigated the performance of contrast and classification-based visual features, emphasizing that DMs share common low-level features, which makes them easy to distinguish. Moreover, They proposed a supervised contrast learning-based method to disentangle semantic and perceptual information, which addresses a more challenging scenario. Zhong et al. [89] focused on the inter-pixel correlation contrast between rich and poor texture regions within an image, and presented a universal detector which can generalize to various AI models, including GAN-based and DM-based models. Existing state-of-the-art detectors have a generalization of cross architectures, but the generalization of cross concepts is not considered. To this end, Dogoulis et al. [90] proposed a sampling strategy that takes into account the image quality scores of the sampled training data, and can effectively improve the detection performance in the cross-concept setting. Innovatively, Bi et al. [91] explored the invariance of real images, and proposed a method to map real images to a dense subspace in the feature space, while all generative images are projected outside this subspace. In this way, it can effectively address longstanding issues in generative detection, e.g., poor generalization, high training costs, and weak interpretability. Generative text detection: Metric-based detection extracts distinguishable features from the generative text. Early on, GLTR [92] was a tool to assist humans in detecting generated text. It employs a set of baseline statistical methods that can detect generation artifacts in common sampling schemes. In a human subject study, the annotation scheme provided by GLTR improved human detection of fake text from 54% to 72% without any prior training. Mitchell et al. [93] noticed that the texts sampled from LLMs tend to occupy negative curvature regions of the model’s log probability function. Based on this, DetectGPT is proposed to set a new curvature-based criterion for detection without additional training. Tulchinskii et al. [94] proposed a new distinguishable representation, the intrinsic dimension. Fluent texts in natural languages have an average intrinsic dimension of 9 or 7 in each language, while AI-generated texts have a lower average intrinsic dimension of 1.5 in each language. Detectors constructed on the basis of intrinsic dimensionality have strong generalizability to models and scenarios. Regarding the model-based methods [95, 96], a classification model is usually trained using a corpus. Guo et al. [95] proposed a text detector for ChatGPT. The detector is based on the RoBERTa model, which is trained by plain answer text and question-answer text pairs respectively. Chen et al. [96] trained two different text classification models using robustly optimized BERT pretraining approach (RoBERTa) and text-to-text Transfer Transformer (T5), respectively, and achieved significant performance on the test dataset with an accuracy of more than 97%.
TABLE IV: A summary of the solutions for detection and attribution of generative data.
Ref. Year Model Category Method
Generative Detection
[83]
2023
DM
Image
Hidden artifacts
[84]
2023
DM
Image
Dual-stream network
[85]
2023
GAN&DM
Image
Inductive bias
[87]
2023
DM
Image
Reconstruction error
[88]
2023
DM
Image
Low-level features
[89]
2023
GAN&DM
Image
Inter-pixel correlation
[90]
2023
GAN&DM
Image
Quality-based sampling
[91]
2023
GAN&DM
Image
Invariance of real images
[92]
2019
GPT-2
Text
Baseline statistical methods
[93]
2023
GPT-3
Text
Curvature-based criterion
[94]
2023
Multi-GPTs
Text
Intrinsic dimension
[95]
2023
Multi-GPTs
Text
Model training
[96]
2023
Multi-GPTs
Text
Model training
Generative Attribution
[97]
2023
Multi-LLMs
Text
Systematic quantification
[98]
2022
GAN
Image
GAN fingerprints
[99]
2023
GAN
Image
Progressive simulation
[100]
2022
DM
Image
Multi-class classifier
[101]
2023
DM
Image
MultiLID
[102]
2023
GAN&DM
Image
Hierarchical multi-level
[86]
2023
DM
Image
Feature retrieval
V-B2 Generative Attribution
He et al. [97] extended current detectors to the potential of text attribution to recognize the source model of a given text. The results show that all these detectors have certain attribution capabilities and still have room for improvement. Moreover, model-based detectors can significantly outperform metric-based detectors. For visual generative data, a lot of attribution works on GANs [103, 98, 104, 105, 99] has been proposed. RepMix [98] is a GAN-fingerprinting technique based on representation mixing and a novel loss. It is able to determine from which structure of GAN a given image is generated. POSE [99] tackles an important challenge, i.e., open-set model attribution, which can simultaneously attribute images to seen and unseen models. POSE simulates open-set samples that keep the same semantics as closed-set samples but embed distinct traces. Recent works have begun to focus on DMs. Sha et al. [100] constructed a multi-class (instead of binary) classifier to attribute fake images generated by DMs. Experiments showed that attributing fake images to their originating models can be achieved effectively, because different models leave unique fingerprints in their generated images. Lorenz et al. [101] designed the multi-local intrinsic dimensionality (multiLID), which is effective in identifying the source diffusion model. Guarnera et al. [102] developed a novel multi-level hierarchical approach based on ResNet models, which can recognize the specific AI architectures (GANs/DMs). The experimental results demonstrate the effectiveness of the proposed approach, with an average accuracy of more than 97%. Intriguingly, a new work [86] can attribute generative data to the training data rather than the source model, necessitating the identification of a subset of training images that contribute most significantly to the generated data. In Fig. 8, we can query the generated data in the training set and evaluate their similarity, which can contribute to protecting the copyright of training data rather than models.
V-C Summary
In table IV, we summarize the solutions for the detection and attribution of generative data. Generative data contains traces (referred to as fingerprints) left by generative models, allowing researchers to detect and attribute the data based on these fingerprints. However, as generative models undergo iterative optimization, these fingerprints also continuously evolve. Therefore, new detection methods must be updated in real time. In contrast, real data exhibits certain invariable features that remain unaltered over time or under varying circumstances. Future detection methods should be designed to effectively harness these invariable features, thereby enhancing the accuracy and robustness of detecting generated data. Furthermore, watermark-based methods have demonstrated notable potential in enhancing tasks related to detection and attribution. However, it’s important to note that such methods constitute an active defense strategy, necessitating preprocessing during data generation. In real-world scenarios, constraining adversaries to exclusively employ watermarked generative data may be not feasible.
VI Compaliance on Generative Data
VI-A Requirements to Compaliance
The compliance of generative data refers to the requirement that such data must adhere to applicable laws, regulations, and ethical standards. With the rapid development and extensive utilization of AIGC technology, the compliance of generative data has become an important topic, encompassing various aspects, e.g., ethics, bias, and politics. Countries and organizations around the world have initiated investigations and issued relevant policies and regulations regarding the regulation of generative data. The United States’ Blueprint for an AI Bill of Rights [106] emphasizes generative data to ensure fairness, privacy protection, and accountability. European Parliament passed the Artificial Intelligence Act [107], which supplements the regulatory regime for generative models and requires that all generative data should be disclosed as derived from AI. China proposed an AIGC-specific regulation, i.e., the Interim Regulation on the Management of Generative Artificial Intelligence Services [12]. This regulation encourages innovation of AIGC but requires a prohibition on the generation of toxicity, e.g., violence, bias, and obscene pornography, as well as an increase in the factuality of the generative data to avoid misleading the public.
Figure 9: Examples of generative images with toxicity from [108].
VI-A1 Non-toxicity
Toxicity presents in generative data involves incongruence with human values or bias directed at particular groups, which has the potential to endanger societal cohesion and intensify divisions among different groups. Since the training of AIGC models is based on a large amount of unintervened data, toxicity in training data [109] directly leads to the corresponding toxicity in generative data, which covers a variety of topics, e.g., sexuality, hatred, politicalization, violentism, and race bias. Some toxic generative examples are shown in Fig. 9. Some works [110, 111] have found stronger associations between males and occupations in language models, verifying the gender bias in generative data. The investigations revealed that GPT-3 consistently and strongly exhibits biased views against the Muslim community [112]. Stable Diffusion v1 was trained on the LAION-2B dataset, which contains images described only in English, making generative data biased towards white culture [113]. Likewise, it was observed that DALLA-E displayed unfavorable biases towards minority groups [114]. Qu et. al [108] provided a comprehensive safety assessment concerning the generation of toxic images, particularly hateful memes from diffusion models. To quantitatively assess the safety of generative images, a safety classifier is developed to identify toxic images based on the predetermined criteria for unsafe content. Their findings indicated that the utilization of harmful prompts resulted in diffusion models producing a significant quantity of toxic images. Additionally, even when prompts are innocuous, the potential for generating toxic images persists. Overall, the danger of large-scale generation of toxic images is imminent.
VI-A2 Factuality
AI-generated data may be contrary to the facts, thus misleading the public. For example, ChatGPT may produce responses that sound reasonable and authoritative but are factually incorrect or nonsensical. Even worse, AIGC often explains its generated responses. When AIGC fails to provide accurate responses to the queries, it not only delivers incorrect information but also supplements seemingly plausible explanations. This enhances the inclination of users to place greater trust in these erroneous contents. The United States news credibility assessment and research organization, NewsGuard, conducted a test on ChatGPT [115]. Researchers posed questions to ChatGPT containing conspiracy theories and misleading narratives and found that it could adapt information within seconds, generating a substantial amount of persuasive yet unattributed content. When used in important domains, such unfactual generative data will bring serious negative impacts [116]. In the healthcare domain, medical diagnosis requires interpretable and correct information. Once AI-generated diagnostic advice is factually incorrect, it will cause irreparable harm to the patient’s life and health. In the journalism domain, news that distorts the facts will mislead the public and undermine the credibility of the media. In the education domain, the dissemination of incorrect knowledge to students will confuse their minds, thus seriously hampering their academic growth and cognitive development.
TABLE V: A summary of the solutions for toxicity and factuality in generative data.
Ref. Year Method Brief introduction
Non-toxicity
[117]
2022
Dataset filtering
Employing the law and legal data to inform data filtering practices.
[118]
2022
Generation guidance
Forgetting the harmful outputs in a confrontational manner.
[119]
2023
Generation guidance
Learned toxic representations for inappropriate mitigation.
[120]
2023
Generation guidance
Extension the generative process by confronting toxic concept.
[121]
2023
Model fine-tuning
Appropriate style to guide the ablation of toxic concept.
[122]
2023
Model fine-tuning
A continual learning-based method to selectively forget concepts.
[123]
2022
Output filtering
Reverse engineer the safety filter and invert toxic embeddings.
Factuality
[124]
2021
Truthfulness standards
Standards definitions and potential ways for AIGC truthfulness.
[125]
2019
Model-based metric
A model-based metric for evaluating the factuality of generated text
[126]
2022
Factual-nucleus sampling
New test set and metrics for factuality enhancement.
[127]
2022
Three-dimensional metric
Sample-level metrics for evaluating faithfulness of generative data.
[128]
2023
Activation classifier
Utilizing the hidden layer activation to discriminate the factuality.
[129]
2023
Multiagent Debate
Multiple models conduct multiple debates to unify the results.
[130]
2023
Feedback learning
Fix the generated data based on the feedback from the tool.
VI-B Countermeasures to Compaliance
VI-B1 Countermeasures to Non-toxicity
Efforts to eliminate toxicity can be divided into four categories. The first one is dataset filtering. A non-toxic training dataset is key to ensuring the security of generative data. Some works [131, 28, 117] have implemented comprehensive processes to filter data contained toxic. OpenAI ensures that any violent or sexual content is removed from DALLA-E2 by carefully filtering [28]. Henderson et. al [117] demonstrated how to extract implicit sanitization rules from the Pile of Law, providing researchers with a pathway to develop more sophisticated data filtering mechanisms. However, large-scale dataset filtering also has unexpected side effects on the downstream performance [132]. The second one is generation guidance. Ganguli et al. [118] identified and attempted to reduce the potentially harmful output of language models in a confrontational manner. They found that the reinforcement learning from human feedback is increasingly difficult to red team as they scale, and a flat trend with scale for the other model types. Brack et al. [119] investigated to instruct effectively diffusion models to suppress inappropriate content using the learned knowledge obtained about the world’s ugliness, thus producing safer and more socially responsible content. Similarly, safe latent diffusion (SLD) [120] extends the generative process via utilizing toxic prompts to guide the safe generation in an opposing direction. The third one is model fine-tuning. Recently, a new term called the ablation concept or ablation forgetting [121, 36, 122] has brought a novel direction to the elimination of toxic content in generative data. Gandikota et. al [121] studied the erasure of toxic concepts from diffusion model weights via model fine-tuning. The proposed method utilizes an appropriate style as a teacher to guide the ablation of the toxic concepts, e.g., sexuality and copyright infringement. Selective Amnesia [122] is a generalized continuous learning framework for concept ablation that applies to different model types and conditional scenarios. It also allows for controlled ablation of concepts that can be specified by the user. However, the ablation concept’s ability to explain the various definitions of toxic concepts remains limited. Lastly, filtering the generated results is also a viable option. Stable Diffusion includes an after-the-fact safety filter [133] to block toxic images. Unfortunately, the filter easily blocks any generated image that is too close to at least one of the 17 predefined sensitive concepts. Rando et. al [123] reverse-engineered this filter and then proposed a manual strategy that enables content not related to sensitive concepts to bypass the filter. In addition, existing toxicity detectors [134, 135] for real data may be able to be updated to be compatible with generative data.
VI-B2 Countermeasures to Factuality
[…]

### v3

@@ 5.2.1. Generative Detection
Generative visual detection: The presence of artifacts in generative images is an important detection cue, which may derive from defects in the generation process or from a specific generative architecture. Corvi et al. (Corvi et al., 2023) gave a preliminary trial to the problem of detecting generative images produced by DMs. Their study showed that the hidden artifact features of images of DMs are partially similar to those observed in images of GANs. Both GANs and DMs leave artifacts on generative data, but only the artifacts are different. Due to the instability of the adversarial training process between the generator and the discriminator, GANs generate unnatural artifacts, e.g., blurred edges and inconsistent textures. While DMs retain small noise features during the stepwise denoising process, which can lead to more natural noise or detail distortion. Xi et al. (Xi et al., 2023) developed a robust dual-stream network consisting of a residual stream and a content stream to capture generic anomalies generated by AIGC. The residual stream utilizes the spatial rich model (SRM) to extract various texture information from images, while the content stream captures additional artifact traces at low frequencies, thus supplementing the residual stream with information that may have been missed. Sinitsa et al. (Sinitsa and Fried, 2024) presented a rule-based method that can achieve high detection accuracy by training a small number of generative images (less than 512). The method employs the inductive bias of CNNs to extract fingerprints of different generators from the training set and applies it to detect generative images of the same model and its fine-tuned versions. Joslin et al. (Joslin et al., 2024) introduced human factors to enhance generative detection. Detecting AI-synthesized faces by combining attentional learning methods with user annotations. They also created a crowdsourcing annotation method to systematically gather various user annotations to identify suspicious areas and extract artifact patterns. With the performance optimization that comes from a larger parametric model, the generative data will more closely mimic the original data while being able to circumvent artifacts, which increases the difficulty of generative detection.
Analyzing distinctive features of generated images is also a viable approach to consider. Interestingly, Wang et al. (Wang et al., 2023a) observed that the image generated by the DMs can be reconstructed by approximating the source model, while the real image cannot. Therefore, they proposed a novel image representation called diffusion reconstruction error (DIRE), which measures the distance between the input images and the reconstructed one. DIRE provides a reliable, simple, and generalized method to differentiate between real images and diffusion-generated images. SeDID (Ma et al., 2023a) leverage the unique properties of diffusion models, namely deterministic inverse and deterministic denoising computational error. In addition, its use of insights from member inference attacks to emphasize distributional differences between real and generative data enhances the understanding of the security and privacy implications of diffusion models. Zhong et al. (Zhong et al., 2023) focused on the inter-pixel correlation contrast between rich and poor texture regions within an image, and presented a universal detector which can generalize to various AI models, including GAN-based and DM-based models. Existing state-of-the-art detectors have a generalization of cross architectures, but the generalization of cross concepts is not considered. To this end, Dogoulis et al. (Dogoulis et al., 2023) proposed a sampling strategy that takes into account the image quality scores of the sampled training data, and can effectively improve the detection performance in the cross-concept setting.
Innovatively, Bi et al. (Bi et al., 2023) explored the invariance of real images, and proposed a method to map real images to a dense subspace in the feature space, while all generative images are projected outside this subspace. In this way, it can effectively address longstanding issues in generative detection, e.g., poor generalization, high training costs, and weak interpretability.
Generative text detection: Metric-based detection extracts distinguishable features from the generative text. Early on, GLTR (Gehrmann et al., 2019) was a tool to assist humans in detecting generated text. It employs a set of baseline statistical methods that can detect generation artifacts in common sampling schemes. In a human subject study, the annotation scheme provided by GLTR improved human detection of fake text from 54% to 72% without any prior training. Mitchell et al. (Mitchell et al., 2023) noticed that the texts sampled from LLMs tend to occupy negative curvature regions of the model’s log probability function. Based on this, DetectGPT is proposed to set a new curvature-based criterion for detection without additional training. Tulchinskii et al. (Tulchinskii et al., 2024) proposed a new distinguishable representation, the intrinsic dimension. Fluent texts in natural languages have an average intrinsic dimension of 9 or 7 in each language, while AI-generated texts have a lower average intrinsic dimension of 1.5 in each language. Detectors constructed on the basis of intrinsic dimensionality have strong generalizability to models and scenarios.
Regarding the model-based methods (Guo et al., 2023; Chen et al., 2023a), a classification model is usually trained using a corpus. Guo et al. (Guo et al., 2023) proposed a text detector for ChatGPT. The detector is based on the RoBERTa model, which is trained by plain answer text and question-answer text pairs respectively. Chen et al. (Chen et al., 2023a) trained two different text classification models using robustly optimized BERT pretraining approach (RoBERTa) and text-to-text Transfer Transformer (T5), respectively, and achieved significant performance on the test dataset with an accuracy of more than 97%.
Table 4. A summary of the representative solutions for the authenticity of generative data.
Ref. Year Remarks Limitations
Genera-tive Detection (Corvi et al., 2023) 2023 Pioneering detection for DMs Lack of JPEG robustness
(Sinitsa and Fried, 2024) 2024 Low budget, multi-model applicable Lack of blurring robustness
(Wang et al., 2023a) 2023 Strong interpretability High resource consumption
(Zhong et al., 2023) 2023 Universal to DMs and GANs Poor results on special models
(Ma et al., 2023a) 2023 Use of DM distinct property Not applicable to GANs
(Dogoulis et al., 2023) 2023 Generalizing cross-conceptes Not high detection performance
(Bi et al., 2023) 2023 High generalization, low costs Imperfect JPEG robustness
(Gehrmann et al., 2019) 2019 Training-free, numerically calculable Unsatisfactory detection accuracy
(Mitchell et al., 2023) 2023 Human-readable, high-efficiency Strong white-box assumption
(Tulchinskii et al., 2024) 2024 Generalization and robustness Failure to small-sample languages
(Guo et al., 2023) 2023 Large-scale data, human evaluations Resource-intensive, poorly explained
(Chen et al., 2023a) 2023 High accuracy, explicable Not scalable, only for English
Genera-tive Attribution (He et al., 2024) 2024 Systematic quantification Medium attribution performance
(Bui et al., 2022) 2022 Robust and practical attribution Only ofr GANs, noise-irrobustness
(Yang et al., 2023) 2023 Open-Set model attribution Moderate versatility and scalability
(Sha et al., 2023) 2023 Pioneering attribution for DMs Lack of generalizability
(lor, 2023) 2023 Lightweight, superior performance Difficult to scale to other models
(Guarnera et al., 2024) 2024 Hierarchical multi-level Unrobust in compression and scaling
(Wang et al., 2023b) 2023 Paradigm of data attribution Lack of strict proof
(Asnani et al., 2024) 2024 Paradigm of concept attribution Requires predefined concepts
5.2.2. Generative Attribution
He et al. (He et al., 2024) extended current detectors to the potential of text attribution to recognize the source model of a given text. The results show that all these detectors have certain attribution capabilities and still have room for improvement. Moreover, model-based detectors can significantly outperform metric-based detectors.
For visual generative data, a lot of attribution works on GANs (Yu et al., 2019; Bui et al., 2022; Yang et al., 2022; Girish et al., 2021; Yang et al., 2023) has been proposed. RepMix (Bui et al., 2022) is a GAN-fingerprinting technique based on representation mixing and a novel loss. It is able to determine from which structure of GAN a given image is generated. POSE (Yang et al., 2023) tackles an important challenge, i.e., open-set model attribution, which can simultaneously attribute images to seen and unseen models. POSE simulates open-set samples that keep the same semantics as closed-set samples but embed distinct traces.
Figure 8. Generative data attributes to the training data from (Wang et al., 2023b).
Recent works have begun to focus on DMs. Sha et al. (Sha et al., 2023) constructed a multi-class (instead of binary) classifier to attribute fake images generated by DMs. Experiments showed that attributing fake images to their originating models can be achieved effectively, because different models leave unique fingerprints in their generated images. Lorenz et al. (lor, 2023) designed the multi-local intrinsic dimensionality (multiLID), which is effective in identifying the source diffusion model. Guarnera et al. (Guarnera et al., 2024) developed a novel multi-level hierarchical approach based on ResNet models, which can recognize the specific AI architectures (GANs/DMs). The experimental results demonstrate the effectiveness of the proposed approach, with an average accuracy of more than 97%.
Intriguingly, a new work (Wang et al., 2023b) can attribute generative data to the training data rather than the source model, necessitating the identification of a subset of training images that contribute most significantly to the generated data. In Fig. 8, we can query the generated data in the training set and evaluate their similarity, which can contribute to protecting the copyright of training data rather than models.
5.3. Summary
In table 4, we summarize the solutions for the authenticity of generative data. Generative data contains traces (referred to as fingerprints) left by generative models, allowing researchers to detect and attribute the data based on these fingerprints. However, as generative models undergo iterative optimization, these fingerprints also continuously evolve. Therefore, new detection methods must be updated in real time. In contrast, real data exhibits certain invariable features that remain unaltered over time or under varying circumstances. Future detection methods should be designed to effectively harness these invariable features, thereby enhancing the accuracy and robustness of detecting generated data. Furthermore, watermark-based methods have demonstrated notable potential in enhancing tasks related to detection and attribution. However, it’s important to note that such methods constitute an active defense strategy, necessitating preprocessing during data generation. In real-world scenarios, constraining adversaries to exclusively employ watermarked generative data may be not feasible. In addition, adversarial attacks can also have an impact on the controllability of generative data. On the one hand, adversarial attacks can attack generative models to prevent them from generating inauthentic content, which ensures the authenticity of generative data at the source. On the other hand, adversarial attacks can attack generative detectors to obfuscate their decisions, which exacerbates the authenticity challenge on generative data.
6. Compaliance on Generative Data
6.1. Requirements to Compaliance
The compliance of generative data refers to the requirement that such data must adhere to applicable laws, regulations, and ethical standards. With the rapid development and extensive utilization of AIGC technology, the compliance of generative data has become an important topic, encompassing various aspects, e.g., ethics, bias, and politics.
Countries and organizations around the world have initiated investigations and issued relevant policies and regulations regarding the regulation of generative data. The United States’ Blueprint for an AI Bill of Rights emphasizes generative data to ensure fairness, privacy protection, and accountability. European Parliament passed the Artificial Intelligence Act, which supplements the regulatory regime for generative models and requires that all generative data should be disclosed as derived from AI. China proposed an AIGC-specific regulation, i.e., the Interim Regulation on the Management of Generative Artificial Intelligence Services (int, 2023). This regulation encourages innovation of AIGC but requires a prohibition on the generation of toxicity, e.g., violence, bias, and obscene pornography, as well as an increase in the factuality of the generative data to avoid misleading the public.
Figure 9. Examples of generative images with toxicity from (Qu et al., 2023).
6.1.1. Non-toxicity
Toxicity presents in generative data involves incongruence with human values or bias directed at particular groups, which has the potential to harm societal cohesion and intensify divisions among different groups. Since the training of AIGC models is based on a large amount of unintervened data, toxicity in training data (Birhane et al., 2021) directly leads to the corresponding toxicity in generative data, which covers a variety of harmful topics, e.g., sexuality, hatred (Birhane et al., 2024), politicalization, violentism, and race bias. Some toxic generative examples are shown in Fig. 9.
Some works (Caliskan et al., 2017; Sheng et al., 2019) have found stronger associations between males and occupations in language models, verifying the gender bias in generative data. The investigations revealed that GPT-3 consistently and strongly exhibits biased views against the Muslim community (Abid et al., 2021). Stable Diffusion v1 was trained on the LAION-2B dataset, which contains images described only in English, making generative data biased towards white culture . Likewise, it was observed that DALLA-E displayed unfavorable biases towards minority groups.
Unlike GANs that only use random noise to generate data, DMs can be guided by additional textual prompts, which increases the risk of generating non-compliant content. As a result, researchers primarily focus on the compliance of data generated by DMs. Qu et. al (Qu et al., 2023) provided a comprehensive safety assessment concerning the generation of toxic images, particularly hateful memes from diffusion models. To quantitatively assess the safety of generative images, a safety classifier is developed to identify toxic images based on the predetermined criteria for unsafe content. Their findings indicated that the utilization of harmful prompts resulted in diffusion models producing a significant quantity of toxic images. Additionally, even when prompts are innocuous, the potential for generating toxic images persists. Overall, the danger of large-scale generation of toxic images is imminent.
6.1.2. Factuality
AI-generated data may be contrary to the facts (Wang et al., 2023c), which is harmful to the public through misleading cognition. For example, ChatGPT may produce responses that sound reasonable and authoritative but are factually incorrect or nonsensical. Even worse, AIGC often explains its generated responses. When AIGC fails to provide accurate responses to the queries, it not only delivers incorrect information but also supplements seemingly plausible explanations. This enhances the inclination of users to place greater trust in these erroneous contents. The United States news credibility assessment and research organization, NewsGuard, conducted a test on ChatGPT (nyt, 2023). Researchers posed questions to ChatGPT containing conspiracy theories and misleading narratives and found that it could adapt information within seconds, generating a substantial amount of persuasive yet unattributed content.
When used in important domains, such unfactual generative data will bring serious harmful impacts (Bender et al., 2021). In the healthcare domain, medical diagnosis requires interpretable and correct information. Once AI-generated diagnostic advice is factually incorrect, it will cause irreparable harm to the patient’s life and health. In the journalism domain, news that distorts the facts will mislead the public and undermine the credibility of the media. In the education domain, the dissemination of incorrect knowledge to students will confuse their minds, thus seriously hampering their academic growth and cognitive development.
6.2. Countermeasures to Compaliance
6.2.1. Countermeasures to Non-toxicity
Efforts to eliminate toxicity can be divided into four categories. The first one is dataset filtering. A non-toxic training dataset is key to ensuring the security of generative data. Some works (Schuhmann et al., 2022; DAL, 2022; Henderson et al., 2022) have implemented comprehensive processes to filter data contained toxic. OpenAI ensures that any violent or sexual content is removed from DALLA-E2 by carefully filtering (DAL, 2022). Henderson et. al (Henderson et al., 2022) demonstrated how to extract implicit sanitization rules from the Pile of Law, providing researchers with a pathway to develop more sophisticated data filtering mechanisms. However, large-scale dataset filtering also has unexpected side effects on the downstream performance (Nichol et al., 2021).
The second one is generation guidance. Ganguli et al. (Ganguli et al., 2022) identified and attempted to reduce the potentially harmful output of language models in a confrontational manner. They found that the reinforcement learning from human feedback is increasingly difficult to red team as they scale, and a flat trend with scale for the other model types. Brack et al. (Brack et al., 2023) investigated to instruct effectively diffusion models to suppress inappropriate content using the learned knowledge obtained about the world’s ugliness, thus producing safer and more socially responsible content. Similarly, safe latent diffusion (SLD) (Schramowski et al., 2023) extends the generative process via utilizing toxic prompts to guide the safe generation in an opposing direction.
The third one is model fine-tuning. Recently, a new term called the ablation concept or ablation forgetting (Gandikota et al., 2023; Kumari et al., 2023; Heng and Soh, 2024) has brought a novel direction to the elimination of toxic content in generative data. Gandikota et. al (Gandikota et al., 2023) studied the erasure of toxic concepts from diffusion model weights via model fine-tuning. The proposed method utilizes an appropriate style as a teacher to guide the ablation of the toxic concepts, e.g., sexuality and copyright infringement. Selective Amnesia (Heng and Soh, 2024) is a generalized continuous learning framework for concept ablation that applies to different model types and conditional scenarios. It also allows for controlled ablation of concepts that can be specified by the user. However, the ablation concept’s ability to explain the various definitions of toxic concepts remains limited. Inspired with social psychological principles, Xu et al. (Xu et al., 2024b) proposed a novel strategy to motivate LLM to integrate different human perspectives and self-regulate their responses.
Lastly, filtering the generated results is also a viable option. Stable Diffusion includes an after-the-fact safety filter to block toxic images. Unfortunately, the filter easily blocks any generated image that is too close to at least one of the 17 predefined sensitive concepts. Rando et. al (Rando et al., 2022) reverse-engineered this filter and then proposed a manual strategy that enables content not related to sensitive concepts to bypass the filter. In addition, existing toxicity detectors (Markov et al., 2023; Lu et al., 2023c) for real data may be able to be updated to be compatible with generative data. This self-correcting mechanism can significantly reduce toxicity and bias.
6.2.2. Countermeasures to Factuality
In order to constrain the non-factualness caused by AIGC “lying”, Evans et al. (Evans et al., 2021) first identified clear standards for AI truthfulness and explored potential ways to establish them.
Table 5. A summary of the representative solutions for the compliance of generative data.
Ref. Year Remarks Limitations
Non-toxicity (Henderson et al., 2022) 2022 Legally guaranteed Only for U.S. texts, perturbation-sensitive
(Ganguli et al., 2022) 2022 The idea of constant confrontation Time-consuming, difficult to scale
(Rando et al., 2022) 2022 Comprehensive toxicity concept Takes a lot of labor, subjective definition
(Brack et al., 2023) 2023 World knowledge guidance Need to modify the model
(Schramowski et al., 2023) 2023 Free-train, maintain quality Lack of provability and ethical concern
(Heng and Soh, 2024) 2024 Controllable forgetting Computing costly, non-automated
(Xu et al., 2024b) 2024 Self-regulate, low cost Ignorance of sensitive vocabularies
Factuality (Lee et al., 2022) 2022 Open-ended test, new metrics Lack of moral consideration
(Alaa et al., 2022) 2022 Three-dimensional metric Divergence measures collapse
(Azaria and Mitchell, 2023) 2023 Simple and powerful Difficult to interpret
(Du et al., 2023) 2023 Comprehensible, reasonable Large resource consumption
(Gou et al., 2024) 2024 Practical and simple Limited to syntactic correctness
(Chen et al., 2024a) 2024 Without extra models or human Difficulty in controlling mistake
(Zhang et al., 2024a) 2024 Learned refusal ability Lack of rigorous evaluation
A reasonable assessment of content factuality is the critical step toward responsible generative data. Goodrich et al. (Goodrich et al., 2019) constructed relation classifiers and fact extraction models based on Wikipedia and Wikidata, by which the factual accuracy of generated text can be measured. Lee et al. (Lee et al., 2022) proposed a novel training method to enhance factuality by utilizing TOPIC PREFIX for better perception of facts and sentence completion as the training objective, which can significantly decrease the number of counterfactuals. They also study the factual accuracy of LLMs with parameter sizes ranging from 126M to 530B and find that the larger the model, the higher the accuracy. This is because the model has a large enough capacity to learn more generalized knowledge, thus reducing the occurrence of counterfactuals.
Alaa et al. (Alaa et al., 2022) designed a three-dimensional metric capable of characterizing the fidelity, diversity, and generalization of generative data from widely generative models in a wide range of applications. SAPLMA (Azaria and Mitchell, 2023) is a simple but powerful method, which only uses the hidden layer activation of LLM to discriminate the factuality of generated statements.
Interestingly, Du et al. (Du et al., 2023) prompted multiple language models to debate their viewpoints and reasoning processes over multiple rounds, and finally come up with a unitive answer. The results indicate that it enhances mathematical and strategic reasoning in the task while reducing the fallacious answers and illusions that modern models are prone to. CRITIC (Gou et al., 2024) requires LLMs to interact with appropriate tools for feedback learning, such as using a search engine for fact-checking or a code interpreter for debugging. The output of the LLM is modified incrementally by the evaluation results of the feedback by the tools. Chen et al. (Chen et al., 2024a) presented a new alignment framework designed to improve LLMs by converting flawed instruction-response pairs into useful alignment data through mistake analysis. By generating harmful responses and analyzing their own mistakes, the framework can improve the alignment with human values. To prevent LLMs from rambling without knowing the answer, Zhang et al. (Zhang et al., 2024a) taught LLMs the ability to refuse to answer. Specifically, they constructed datasets of refusal perception, and then adapted the model to avoid answering questions that were beyond its parametric knowledge.
6.3. Summary
In table 5, we summarize the solutions for toxicity and factuality in generative data. The first problem to be solved for the compliance of generative data is to define its standards. In addition, there should be different compliance standards for different application scenarios, rather than a blanket denial of generative content. For example, “a flying pig” may not be factual for a language model, but it is more creative for a visual model. After that, the generated content of AI can be made compliant with the specification within the constraints of the standard. In addition, adversarial attacks can also have an impact on the controllability of generative data. On the one hand, adversarial attacks can attack the generative model to prevent it from learning the non-compliant content in the real data, which enhances the compliance of generative data. On the other hand, adversarial attacks can also attack compliance detectors to evade their detection, which exacerbates the compliance challenge for generative data.
In conclusion, we emphasize the examination of toxicity and factuality issues within the context of compliance to ensure the responsible and ethical use of AI-generated content across various domains. Nevertheless, existing approaches exhibit certain limitations, nec essitating continuous efforts by researchers to address these shortcomings and better align with the practical requirements of real-world applications.
7. Benchmark and Statistical Analysis
7.1. Benchmark
7.1.1. Benchmark Dataset
[…]